A company had stock outstanding as follows during each of its first three years of operations:...

  1. A company had stock outstanding as follows during each of its first three years of operations: 2,000 shares of 8%, $100 par, cumulative preferred stock and 34,000 shares of $10 par common stock. The amounts distributed as dividends are presented below. Determine the total and per-share dividends for each class of stock for each year by completing the schedule. Round dividends per share to the nearest cent. Enter "0" if no dividends are paid.

    Preferred

    Common

    Year

    Dividends

    Total

    Per Share

    Total

    Per Share

    1 $12,000   $ $ $ $
    2 16,000 $ $ $ $
    3 24,740 $ $ $ $
Answer #1
Preferred Common
Year Dividends Total Per Share Total Per Share
1        12,000.0 12,000 12,000/2,000 = 6 0 0  
2        16,000.0 16,000 16,000/2,000 = 8 0   0
3        24,740.0 20,000 20,000/2,000 = 10 4,740.00 4,740/34,000 = 0.13

Annual preferred dividend = Number of preferred share x par value per share x Dividend rate

= 2,000 x 100 x %

= $16,000

Preferred dividend in arrears in year 1 = $4,000

In the third year, dividend paid to preferred stock holders = 16,000+4,000

= $20,000

Kindly comment if you need further assistance. Thanks‼!